# Geometry and factorization of multivariate Markov chains with applications to the swapping algorithm

### v4

@@ 2.1 Distance to independence and the closest product chain
$\mathbb{I}_{f}^{\pi}(P):=\min_{L_{i}\in\mathcal{L}(\mathcal{X}^{(i)}),\,\forall i\in\llbracket d\rrbracket}D_{f}^{\pi}(P\|\otimes_{i=1}^{d}L_{i}).$ (4)
@@
Note that since $L\mapsto D_{f}^{\pi}(M\|L)$ is continuous and the set ×$\bigtimes_{i=1}^{d}\mathcal{L}(\mathcal{X}^{(i)})$ is compact, the minimization problem (4) is always attained. The next result states that this distance to independence of $P$ is zero if and only if $P$ is a product chain under suitable assumptions on $\pi$. This is analogous to the property that if two random variables are independent, then their correlation is zero.
@@
Proof.
The non-negativity is clear from Proposition 2.1. If $P$ is a product chain, then clearly $\mathbb{I}^{\pi}_{f}(P)=0$. For the other direction, if $\mathbb{I}^{\pi}_{f}(P)=0$, then $D_{f}^{\pi}(P\|\otimes_{i=1}^{d}L_{i})=0$ for some $L_{i}$ since the minimization in (4) is exactly attained. By Proposition 2.1, $P=\otimes_{i=1}^{d}L_{i}$. ∎
Let us now recall the notion of edge measure of a Markov chain (see e.g. (Levin and Peres, 2017, equation (7.5))). Let $\pi\in\mathcal{P}(\mathcal{X})$ and $P\in\mathcal{L}(\mathcal{X})$. The edge measure $\pi\boxtimes P\in\mathcal{P}(\mathcal{X}\times\mathcal{X})$ is defined to be, for $x,y\in\mathcal{X}$,
@@
First, we note that $P^{(i)}_{\pi}$ can be understood as a special case of the keep-$S$-in transition matrix to be introduced in Definition 2.4 below, which is a further special case of various notions of “projection chains” investigated in Jerrum et al. (2004); Aldous and Fill (2002); Chen et al. (1999). Second, we see that if $P$ is $\pi$-stationary, then $P^{(i)}_{\pi}$ is $\pi^{(i)}$-stationary. Similarly, if $P$ is $\pi$-reversible, then $P^{(i)}_{\pi}$ is $\pi^{(i)}$-reversible. As a result, $\otimes_{i=1}^{d}P^{(i)}_{\pi}$ is thus $\otimes_{i=1}^{d}\pi^{(i)}$-stationary, and hence in general $\otimes_{i=1}^{d}P^{(i)}_{\pi}$ is not $\pi$-stationary. In the case where $\pi=\otimes_{i=1}^{d}\pi^{(i)}$ is a product stationary distribution, then $\otimes_{i=1}^{d}P^{(i)}_{\pi}$ is $\pi$-stationary. A generalization of the above discussions can be found in Proposition 2.3 below.
@@
Suppose that $P$ is $\pi$-stationary, and $(\mathbf{X}_{n})_{n\in\mathbb{N}}=(X^{1}_{n},X^{2}_{n},\ldots,X^{d}_{n})_{n\in\mathbb{N}}$ (resp. $(\mathbf{Y}_{n})_{n\in\mathbb{N}}=(Y^{1}_{n},Y^{2}_{n},\ldots,Y^{d}_{n})_{n\in\mathbb{N}}$) is the discrete-time homogeneous Markov chain with transition matrix $P$ (resp. $\otimes_{i=1}^{d}P^{(i)}_{\pi}$). Then, the distance to independence of $P$ can be written as
@@
where $\mathbf{X}_{n}\sim\mu_{n},\mathbf{Y}_{n}\sim\nu_{n}$ and the right hand side can be interpreted as the KL divergence rate from the closest product chain to $P$. The rightmost expression in the equality above is also known as the mutual information rate Chen et al. (2018); Holliday et al. (2002).
Proof.
@@
One possible application of the Pythagorean identity lies in assessing the convergence to equilibrium of $P$. This is in part motivated by (Hsu et al., 2019, Section $10$) which suggests looking into “Markov chains with factored transition kernels with a few factors”. Suppose that $P$ is ergodic (i.e. irreducible and aperiodic) with a product form stationary distribution $\pi=\otimes_{i=1}^{d}\pi^{(i)}$. Let $\Pi\in\mathcal{L}(\mathcal{X})$ be the matrix where each row is $\pi$, and $\Pi^{(i)}\in\mathcal{L}(\mathcal{X}^{(i)})$ be a matrix where each row is $\pi^{(i)}$ for all $i$. We thus see that
@@ 2.1.1 The closest product chain with prescribed marginals and a large deviation principle of Markov chains
    $L_{*}^{(i)}(x^{i},y^{i})\propto\prod_{x^{(-i)},y^{(-i)}}P(x,y)^{\frac{\pi(x)\mathbf{Z}(x^{(-i)},y^{(-i)})}{Z(x^{i},y^{i})}},$ (5)
@@
that is, $L_{*}^{(i)}=P^{(i)}_{\pi}$. ∎
One application of Theorem 2.2 lies in the large deviation analysis and Sanov’s theorem of Markov chains, in which we apply the results obtained in Fayolle and de La Fortelle (2002). We refer readers to Natarajan (1985); Vidyasagar (2014); Dembo and Zeitouni (2010) for related literature on large deviations of Markov chains.
@@
$E_{n}:=\dfrac{1}{n}\left(\sum_{i=1}^{n-1}\delta_{(X_{i},X_{i+1})}+\delta_{(X_{n},X_{1})}\right).$ (6)
@@
:=$:=\{(\otimes_{j=1}^{i-1}L_{j})\otimes M\otimes(\otimes_{j=i+1}^{d}L_{j});~{}\textrm{$\pi^{(i)}$-stationary}\quad M\in\mathcal{L}(\mathcal{X}^{(i)})\},$
@@
Proof.
The plan is to invoke Theorem $1.1$ in Fayolle and de La Fortelle (2002). Let us first assume that $K$ is a subset of the set of balanced measures. Then by Fayolle and de La Fortelle (2002) and Theorem 2.2, these yield
@@ 2.1.2 A coordinate descent algorithm for finding the closest product chain
$\mu^{(-S)}(x^{(-S)}):=\sum_{i\in S}\sum_{x^{i}\in\mathcal{X}^{(i)}}\mu(x^{1},\ldots,x^{d}).$
@@
[…]
$\mu^{(S)}(x^{(S)}):=\mu^{(-\llbracket d\rrbracket\backslash S)}(x^{(-\llbracket d\rrbracket\backslash S)})=\sum_{i\notin S}\sum_{x^{i}\in\mathcal{X}^{(i)}}\mu(x^{1},\ldots,x^{d}).$
@@
Consider a sequence $(S_{i})_{i=1}^{n}$ with $S_{i}\subseteq\llbracket d\rrbracket$, where each $j\in\llbracket d\rrbracket$ belongs to at least $r$ of $S_{i}$. The Shearer’s lemma of entropy (Polyanskiy and Wu, 2022, Theorem $1.8$) is given by
$H(X_{1},\ldots,X_{d})\leq\dfrac{1}{r}\sum_{i=1}^{n}H((X_{l})_{l\in S_{i}}),$ (7)
where $(X_{l})_{l\in S_{i}}\sim\mu^{(S_{i})}$ and $H((X_{l})_{l\in S_{i}}):=-\sum_{x^{(S_{i})}}\mu^{(S_{i})}(x^{(S_{i})})\ln\mu^{(S_{i})}(x^{(S_{i})})$ is the entropy of $\mu^{(S_{i})}$, while the KL divergence version of the Shearer’s lemma (Gavinsky et al., 2015, Corollary $2.8$) is stated as, for $\mu,\nu=\otimes_{j=1}^{d}\nu_{j}\in\mathcal{P}(\mathcal{X})$,
$\widetilde{D}_{KL}(\mu\|\otimes_{j=1}^{d}\nu_{j})\geq\dfrac{1}{r}\sum_{i=1}^{n}\widetilde{D}_{KL}(\mu^{(S_{i})}\|\nu^{(S_{i})}).$ (8)
In the special case of taking $S_{i}=\llbracket d\rrbracket\backslash\{i\}$ and $n=d$ so that $r=d-1$, we recover the Han’s inequality for KL divergence between discrete probability masses (Boucheron et al., 2013, Theorem $4.9$):
$\widetilde{D}_{KL}(\mu\|\otimes_{j=1}^{d}\nu_{j})\geq\dfrac{1}{d-1}\sum_{i=1}^{n}\widetilde{D}_{KL}(\mu^{(-i)}\|\nu^{(-i)}).$ (9)
@@
Viewing these projections as conditional expectations is therefore in line with conditional expectations in the drift term of independent projections in Lacker (2023). In addition, this observation allows us to simulate one step of the projection chain associated with $P^{(-S)}_{\pi}$. Suppose that $\pi(\cdot|x^{(-S)})$ can be sampled from. Starting from the initial state $x^{(-S)}$, we first draw a random $x^{(S)}$ according to $\pi(\cdot|x^{(-S)})$, followed by one step of $P$ from $(x^{(S)},x^{(-S)})$ to $(y^{(S)},y^{(-S)})$. This is applied in Section 3 to design a projection sampler for the swapping algorithm.
We note that when $P$ is ergodic and admits $\pi$ as stationary distribution, the keep-$S$-in transition matrix $P^{(S)}$ can be viewed as a special case of the “projection chain” of $P$ in Pillai and Smith (2017); Madras and Randall (2002); Jerrum et al. (2004), or as an “induced chain” of $P$ in (Aldous and Fill, 2002, Section $4.6$). The latter is also known as a “collapsed chain” in Chen et al. (1999), which can be understood as the opposite of the lifting procedure in MCMC.
Let $\Omega_{x^{(S)}}:=\{x^{(S)}\}\times\mathcal{X}^{(-S)}$. We then see that $(\Omega_{x^{(S)}})_{x^{(S)}\in\mathcal{X}^{(S)}}$ is a partition of the state space, that is, $\mathcal{X}=\bigtimes_{i=1}^{d}\mathcal{X}^{(i)}=\cup_{x^{(S)}\in\mathcal{X}^{(S)}}\Omega_{x^{(S)}}$. Let $\overline{P}$ be the “projection chain” of $P$ with respect to $(\Omega_{x^{(S)}})_{x^{(S)}\in\mathcal{X}^{(S)}}$ in the sense of Jerrum et al. (2004), which is defined to be
@@
We summarize the ergodicity properties of $P^{(S)}$ appeared in Aldous and Fill (2002); Jerrum et al. (2004). We can understand that these properties are inherited from the counterpart properties of the original $P$:
@@
In the context of MCMC, there are often situations in which we are only interested in sampling from a subset $S$ out of the $d$ coordinates of the stationary distribution $\pi$ of a sampler. Thus, $P^{(S)}_{\pi}$ offers a natural projection sampler to approximately sample from $\pi^{(S)}$. As a concrete example, we can consider the run-and-tumble Markov chains (van Ginkel et al., 2021) where the algorithm maintains both the positions and directions of $d$ particles. In such setting, we are often interested in sampling from the stationary distribution of the positions of the particles only and discard the samples from the directions. Another concrete example would be the swapping algorithm that we shall discuss in Section 3, where one maintains a system of Markov chains over a range of temperatures. The swapping algorithm is designed to sample from the Boltzmann-Gibbs distribution at the lowest temperature of the algorithm and samples that correspond to higher temperatures are often discarded. A third example is the auxiliary MCMC methods (Higdon, 1998), where one artificially add auxiliary variables in a MCMC algorithm to improve convergence, and at the end of simulation the samples of these auxiliary variables are discarded. In these algorithms, it is therefore natural to consider $P^{(S)}_{\pi}$ as a candidate sampler for $\pi^{(S)}$. We shall compare hitting and mixing time parameters between $P$ and $P^{(S)}_{\pi}$ in Section 2.5. In addition, we implement these ideas and propose an improved projection sampler based on the swapping algorithm and analyze its mixing time in Section 3.1.
Now, let us recall the partition lemma for KL divergence of probability masses (Brémaud, 2017, Lemma $13.1.3$), which is a consequence of the log-sum inequality. For $\mu,\nu\in\mathcal{P}(\mathcal{X})$ and $\emptyset\neq S\subseteq\llbracket d\rrbracket$, the partition lemma gives
@@
where we apply the log-sum inequality. ∎
@@
In the special case of taking $S_{i}=\llbracket d\rrbracket\backslash\{i\}$ and $n=d$ so that $r=d-1$, we obtain a Han’s inequality of the form
@@
In the special case of taking $S_{i}=\llbracket d\rrbracket\backslash\{i\}$ and $n=d$ so that $r=d-1$, we obtain a Han’s inequality of the form
@@
$\underbrace{\mathbb{I}^{\pi}(P)}_{\text{distance to independence of $P$}}$ $=\underbrace{\mathbb{I}^{\pi}(P,\mathcal{L}_{\otimes_{i=1}^{n}S_{i}}(\mathcal{X}))}_{\text{distance to $(S_{i})_{i=1}^{n}$-factorizability of $P$}}+\underbrace{\mathbb{I}^{\pi}(\otimes_{i=1}^{n}P^{(S_{i})}_{\pi})}_{\text{distance to independence of $\otimes_{i=1}^{n}P^{(S_{i})}_{\pi}$}}$
$=\underbrace{\mathbb{I}^{\pi}(P,\mathcal{L}_{\otimes_{i=1}^{n}S_{i}}(\mathcal{X}))}_{\text{distance to $(S_{i})_{i=1}^{n}$-factorizability of $P$}}+\sum_{i=1}^{n}\underbrace{\mathbb{I}^{\pi^{(S_{i})}}(P^{(S_{i})}_{\pi})}_{\text{distance to independence of $P^{(S_{i})}_{\pi}$}}.$
Proof.
@@
$P$$(\mathcal{L}(\mathcal{X}),D^{\pi}_{KL})$$\otimes_{i=1}^{n}P^{(S_{i})}_{\pi}$$\mathcal{L}_{\otimes_{i=1}^{n}S_{i}}(\mathcal{X})$$\otimes_{i=1}^{d}P^{(i)}_{\pi}$$\mathcal{L}_{\otimes_{i=1}^{l}V_{i}}(\mathcal{X})$$\otimes_{i=1}^{l}P^{(V_{i})}_{\pi}$

### v5

@@ 2.1 Distance to independence and the closest product chain
(5) $\mathbb{I}_{f}^{\pi}(P):=\min_{L_{i}\in\mathcal{L}(\mathcal{X}^{(i)}),\,\forall i\in\llbracket d\rrbracket}D_{f}^{\pi}(P\|\otimes_{i=1}^{d}L_{i}).$
@@
Note that since $L\mapsto D_{f}^{\pi}(M\|L)$ is continuous and the set $\bigtimes_{i=1}^{d}\mathcal{L}(\mathcal{X}^{(i)})$ is compact, the minimization problem (5) is always attained. The next result states that this distance to independence of $P$ is zero if and only if $P$ is a product chain under suitable assumptions on $\pi$. This is analogous to the property that if two random variables are independent, then their correlation is zero.
@@
Proof 2.9.
The non-negativity is clear from Proposition 2.3. If $P$ is a product chain, then clearly $\mathbb{I}^{\pi}_{f}(P)=0$. For the other direction, if $\mathbb{I}^{\pi}_{f}(P)=0$, then $D_{f}^{\pi}(P\|\otimes_{i=1}^{d}L_{i})=0$ for some $L_{i}$ since the minimization in (5) is exactly attained. By Proposition 2.3, $P=\otimes_{i=1}^{d}L_{i}$.
Let us now recall the notion of edge measure of a Markov chain (see e.g. [25, equation (7.5)]). Let $\pi\in\mathcal{P}(\mathcal{X})$ and $P\in\mathcal{L}(\mathcal{X})$. The edge measure $\pi\boxtimes P\in\mathcal{P}(\mathcal{X}\times\mathcal{X})$ is defined to be, for $x,y\in\mathcal{X}$,
@@
First, we note that $P^{(i)}_{\pi}$ can be understood as a special case of the keep-$S$-in transition matrix to be introduced in Definition 2.18 below, which is a further special case of various notions of “projection chains” investigated in [23, 1, 5]. Second, we see that if $P$ is $\pi$-stationary, then $P^{(i)}_{\pi}$ is $\pi^{(i)}$-stationary. Similarly, if $P$ is $\pi$-reversible, then $P^{(i)}_{\pi}$ is $\pi^{(i)}$-reversible. As a result, $\otimes_{i=1}^{d}P^{(i)}_{\pi}$ is thus $\otimes_{i=1}^{d}\pi^{(i)}$-stationary, and hence in general $\otimes_{i=1}^{d}P^{(i)}_{\pi}$ is not $\pi$-stationary. In the case where $\pi=\otimes_{i=1}^{d}\pi^{(i)}$ is a product stationary distribution, then $\otimes_{i=1}^{d}P^{(i)}_{\pi}$ is $\pi$-stationary. A generalization of the above discussions can be found in Proposition 2.20 below.
@@
[…]
@@
where $\mathbf{X}_{n}\sim\mu_{n},\mathbf{Y}_{n}\sim\nu_{n}$ and the right hand side can be interpreted as the KL divergence rate from the closest product chain to $P$. The rightmost expression in the equality above is also known as the mutual information rate [7, 21].
Proof 2.13.
@@
One possible application of the Pythagorean identity lies in assessing the convergence to equilibrium of $P$. This is in part motivated by [22, Section $10$] which suggests looking into “Markov chains with factored transition kernels with a few factors”. Suppose that $P$ is ergodic (i.e. irreducible and aperiodic) with a product form stationary distribution $\pi=\otimes_{i=1}^{d}\pi^{(i)}$. Let $\Pi\in\mathcal{L}(\mathcal{X})$ be the matrix where each row is $\pi$, and $\Pi^{(i)}\in\mathcal{L}(\mathcal{X}^{(i)})$ be a matrix where each row is $\pi^{(i)}$ for all $i$. We thus see that
@@ 2.1.1 The closest product chain with prescribed marginals and a large deviation principle of Markov chains
    (6) $L_{*}^{(i)}(x^{i},y^{i})\propto\prod_{x^{(-i)},y^{(-i)}}P(x,y)^{\frac{\pi(x)\mathbf{Z}(x^{(-i)},y^{(-i)})}{Z(x^{i},y^{i})}},$
@@
that is, $L_{*}^{(i)}=P^{(i)}_{\pi}$.
One application of Theorem 2.14 lies in the large deviation analysis and Sanov’s theorem of Markov chains, in which we apply the results obtained in [14]. We refer readers to [34, 41, 10] for related literature on large deviations of Markov chains.
@@
(7) $E_{n}:=\dfrac{1}{n}\left(\sum_{i=1}^{n-1}\delta_{(X_{i},X_{i+1})}+\delta_{(X_{n},X_{1})}\right).$
@@
$:=\{(\otimes_{j=1}^{i-1}L_{j})\otimes M\otimes(\otimes_{j=i+1}^{d}L_{j});~\textrm{$\pi^{(i)}$-stationary}\quad M\in\mathcal{L}(\mathcal{X}^{(i)})\},$
@@
Proof 2.17.
The plan is to invoke Theorem $1.1$ in [14]. Let us first assume that $K$ is a subset of the set of balanced measures. Then by [14] and Theorem 2.14, these yield
@@ 2.1.2 A coordinate descent algorithm for finding the closest product chain
$\mu^{(-S)}(x^{(-S)}):={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{x^{i}\in\mathcal{X}^{(i)}\,\textrm{for each}\,i\in S}}\mu(x^{1},\ldots,x^{d}).$
@@
The keep-$S$-in distribution of $\mu$ is the leave-$\S\llbracket d\rrbracket\backslash$ S-out distribution of $\mu$, that is,
$\mu^{(S)}(x^{(S)}):=\mu^{(-\llbracket d\rrbracket\backslash S)}(x^{(-\llbracket d\rrbracket\backslash S)})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{x^{i}\in\mathcal{X}^{(i)}\,\textrm{for each}\,i\notin S}}\mu(x^{1},\ldots,x^{d}).$
@@
Consider a sequence $(S_{i})_{i=1}^{n}$ with $S_{i}\subseteq\llbracket d\rrbracket$, where each $j\in\llbracket d\rrbracket$ belongs to at least $r$ of $S_{i}$. The Shearer’s lemma of entropy [36, Theorem $1.8$] is given by
(8) $H(X_{1},\ldots,X_{d})\leq\dfrac{1}{r}\sum_{i=1}^{n}H((X_{l})_{l\in S_{i}}),$
where $(X_{l})_{l\in S_{i}}\sim\mu^{(S_{i})}$ and $H((X_{l})_{l\in S_{i}}):=-\sum_{x^{(S_{i})}}\mu^{(S_{i})}(x^{(S_{i})})\ln\mu^{(S_{i})}(x^{(S_{i})})$ is the entropy of $\mu^{(S_{i})}$, while the KL divergence version of the Shearer’s lemma [15, Corollary $2.8$] is stated as, for $\mu,\nu=\otimes_{j=1}^{d}\nu_{j}\in\mathcal{P}(\mathcal{X})$,
(9) $\widetilde{D}_{KL}(\mu\|\otimes_{j=1}^{d}\nu_{j})\geq\dfrac{1}{r}\sum_{i=1}^{n}\widetilde{D}_{KL}(\mu^{(S_{i})}\|\nu^{(S_{i})}).$
In the special case of taking $S_{i}=\llbracket d\rrbracket\backslash\{i\}$ and $n=d$ so that r=d1r=d-1, we recover the Han’s inequality for KL divergence between discrete probability masses [3, Theorem $4.9$]:
(10) $\widetilde{D}_{KL}(\mu\|\otimes_{j=1}^{d}\nu_{j})\geq\dfrac{1}{d-1}\sum_{i=1}^{n}\widetilde{D}_{KL}(\mu^{(-i)}\|\nu^{(-i)}).$
@@
Viewing these projections as conditional expectations is therefore in line with conditional expectations in the drift term of independent projections in [24]. In addition, this observation allows us to simulate one step of the projection chain associated with $P^{(-S)}_{\pi}$. Suppose that $\pi(\cdot|x^{(-S)})$ can be sampled from. Starting from the initial state $x^{(-S)}$, we first draw a random $x^{(S)}$ according to $\pi(\cdot|x^{(-S)})$, followed by one step of $P$ from $(x^{(S)},x^{(-S)})$ to $(y^{(S)},y^{(-S)})$. This is applied in Section 3 to design a projection sampler for the swapping algorithm.
We note that when $P$ is ergodic and admits $\pi$ as stationary distribution, the keep-$S$-in transition matrix $P^{(S)}$ can be viewed as a special case of the “projection chain” of $P$ in [35, 28, 23], or as an “induced chain” of $P$ in [1, Section $4.6$]. The latter is also known as a “collapsed chain” in [5], which can be understood as the opposite of the lifting procedure in MCMC.
Let $\Omega_{x^{(S)}}:=\{x^{(S)}\}\times\mathcal{X}^{(-S)}$. We then see that $(\Omega_{x^{(S)}})_{x^{(S)}\in\mathcal{X}^{(S)}}$ is a partition of the state space, that is, $\mathcal{X}=\bigtimes_{i=1}^{d}\mathcal{X}^{(i)}=\cup_{x^{(S)}\in\mathcal{X}^{(S)}}\Omega_{x^{(S)}}$. Let $\overline{P}$ be the “projection chain” of $P$ with respect to $(\Omega_{x^{(S)}})_{x^{(S)}\in\mathcal{X}^{(S)}}$ in the sense of [23], which is defined to be
@@
We summarize the ergodicity properties of $P^{(S)}$ appeared in [1, 23]. We can understand that these properties are inherited from the counterpart properties of the original $P$:
@@
In the context of MCMC, there are often situations in which we are only interested in sampling from a subset $S$ out of the $d$ coordinates of the stationary distribution $\pi$ of a sampler. Thus, $P^{(S)}_{\pi}$ offers a natural projection sampler to approximately sample from $\pi^{(S)}$. We also recall that in Section 1.1 we have presented numerical evidence to support the use of projection samplers as motivation of this paper.
As a concrete example, we can consider the run-and-tumble Markov chains [40] where the algorithm maintains both the positions and directions of $d$ particles. In such setting, we are often interested in sampling from the stationary distribution of the positions of the particles only and discard the samples from the directions. Another concrete example would be the swapping algorithm that we shall discuss in Section 3, where one maintains a system of Markov chains over a range of temperatures. The swapping algorithm is designed to sample from the Boltzmann-Gibbs distribution at the lowest temperature of the algorithm and samples that correspond to higher temperatures are often discarded. A third example is the auxiliary MCMC methods [19], where one artificially add auxiliary variables in a MCMC algorithm to improve convergence, and at the end of simulation the samples of these auxiliary variables are discarded. In these algorithms, it is therefore natural to consider $P^{(S)}_{\pi}$ as a candidate sampler for $\pi^{(S)}$. We shall compare hitting and mixing time parameters between $P$ and $P^{(S)}_{\pi}$ in Section 2.5. In addition, we implement these ideas and propose an improved projection sampler based on the swapping algorithm and analyze its mixing time in Section 3.1.
Now, let us recall the partition lemma for KL divergence of probability masses [4, Lemma $13.1.3$], which is a consequence of the log-sum inequality. For $\mu,\nu\in\mathcal{P}(\mathcal{X})$ and $\emptyset\neq S\subseteq\llbracket d\rrbracket$, the partition lemma gives
@@
where we apply the log-sum inequality.
@@
In the special case of taking $S_{i}=\llbracket d\rrbracket\backslash\{i\}$ and $n=d$ so that r=d1r=d-1, we obtain a Han’s inequality of the form
@@
In the special case of taking $S_{i}=\llbracket d\rrbracket\backslash\{i\}$ and $n=d$ so that r=d1r=d-1, we obtain a Han’s inequality of the form
@@
[…]
Proof 2.34.
@@
[…]